# Pairs of Clark Unitary Operators on the Bidisk and their Taylor Joint Spectra

### v1

@@ 5. RIF Level Sets and Taylor Joint Spectrum
By Theorem 3.3, the set $\mathcal{C}_{\alpha}$ is the support of $\sigma_{\alpha}$. Now we show that $\sigma_{\alpha}$ satisfies conditions (a) and (b) from Theorem 5.1.
For (a), by way of contradiction, assume that there is some $\lambda\in\mathbb{T}$ such that $\sigma_{\alpha}(\{\lambda\}\times\mathbb{T})\neq 0$. Since $\sigma_{\alpha}$ does not have any point masses, this implies that there is a countable set of points of the form $(\lambda,\tau_{n})$ with $\tau_{n}\in\mathbb{T},n\in\mathbb{N}$ such that $\phi(\lambda,\tau_{n})=\alpha$. Since the function $\phi(\lambda,\cdot)$ is a finite Blaschke product, this implies that $\phi(\lambda,\cdot)$ must be constant and equal to $\alpha$. But, then $(\{\lambda\}\times\mathbb{T})\subseteq\mathcal{C}_{\alpha}$, which contradicts the fact that $\phi$ is generic. Thus, $\sigma_{\alpha}(\{\lambda\}\times\mathbb{T})=0$ for each $\lambda\in\mathbb{T}$ and an analogous argument gives $\sigma_{\alpha}(\mathbb{T}\times\{\lambda\})=0$ as well.
For (b), we just note that $\mathcal{C}_{\alpha}$ is the closure of the set
@@
as needed. Then the final conclusion follows from Theorem 5.1. ∎
Theorem 5.5.
Let $\phi$ be a rational inner function on $\mathbb{D}^{2}$ and let $\alpha\in\mathbb{T}$ be generic for $\phi$. Then there is a commuting pair of unitary perturbations $(U_{\alpha}^{1},U_{\alpha}^{2})$ of the compressed shifts $S_{\phi}^{1},S_{\phi}^{2}$ such that their Taylor joint spectrum $\boldsymbol{\sigma}(U_{\alpha}^{1},U_{\alpha}^{2})=\mathcal{C}_{\alpha}$.
Proof.
By Corollary 4.6, $J_{\alpha}$ is unitary and there are unitary maps $U_{\alpha}^{1},U_{\alpha}^{2}:K_{\phi}\rightarrow K_{\phi}$ such that $U^{j}_{\alpha}=J_{\alpha}^{*}M_{z_{j}}J_{\alpha}$ for $j=1,2$. Since $M_{z_{1}},M_{z_{2}}$ commute on $L^{2}(\mu_{\alpha})$, operators $U_{\alpha}^{1},U_{\alpha}^{2}$ commute on $K_{\phi}$. Moreover,
@@
where $\vartheta^{2}_{\alpha}(z)=\bar{\alpha}\frac{(B_{2}\phi)(z)}{1-\bar{\alpha}\phi(z_{1},0)}$ and $B_{2}$ denotes the backward shift in $z_{2}$. (As an aside, while Corollary 4.6 only states the information about $U_{\alpha}^{1}$, the proof immediately implies the corresponding information about $U_{\alpha}^{2}$.) Thus, the operators $U^{1}_{\alpha},U^{2}_{\alpha}$ are commuting unitary perturbations of the compressed shifts $S_{\phi}^{1},S_{\phi}^{2}$. By Corollary 5.4, $\boldsymbol{\sigma}(M_{z_{1}},M_{z_{2}})=\mathcal{C}_{\alpha}$. Since the Taylor joint spectrum is invariant under unitary conjugation, this gives $\boldsymbol{\sigma}(U_{\alpha}^{1},U_{\alpha}^{2})=\mathcal{C}_{\alpha}$. ∎
@@
There are several techniques for constructing rational inner functions $\phi$ whose level sets $\mathcal{C}_{\alpha}$ have prescribed properties, see for instance [11, Section 7]. By Theorem 5.5, this in turn allows us to prescribe such curve subsets of $\mathbb{T}^{2}$ as Taylor spectra of natural pairs of commuting unitaries.
@@ 6.1. Finite Blaschke Products
Here, we use inner functions constructed from finite Blaschke products to explore our various results. Specifically, let $\phi_{1}$ and $\phi_{2}$ be finite Blaschke products and consider $\phi(z)=\phi(z_{1})\phi(z_{2})$.
@@
First, let us assume that one of $\phi_{1}$ or $\phi_{2}$ is constant, so that $\phi$ is a function of one variable. Then for each $\alpha\in\mathbb{T},$ the level set $\mathcal{C}_{\alpha}$ will contain lines of either the form $z_{1}=\tau$ or $z_{2}=\tau$ for some $\tau\in\mathbb{T}.$ Then by applying the argument in the proof of Proposition 3.10 in [9], one can show that the isometry $J_{\alpha}$ linking $K_{\phi}$ and $L^{2}(\sigma_{\alpha})$ is not unitary. Thus, our machinery breaks down and we cannot apply Theorem 4.3 to obtain a unitary perturbation of $S_{\phi}^{1}.$
@@
For this reason, it makes sense to restrict attention to $\phi_{1},\phi_{2}$ that are both nonconstant. In this situation, Theorem 4.2 in [4] implies that every $\alpha\in\mathbb{T}$ is generic for $\phi$. From this, we can apply Corollary 4.6 to obtain the formula
@@
for the associated unitary perturbation $U_{\alpha}^{1}$ of $S_{\phi}^{1}$.
@@
If $\phi_{1}(0)=0$, then $\psi_{1}:=\phi_{1}/z$ is also a finite Blaschke product and $\phi(0,z_{2})\equiv 0$. In this case, we have a particularly nice formula for $U_{\alpha}^{1}.$ We can apply Corollary 4.5 to conclude that
@@
where $K_{\phi_{2}}$ is invariant under $T_{1},$ the shift in $z_{1}.$ Thus, $S_{\phi}^{1}|_{K_{\phi_{2}}}=T_{1}|_{K_{\phi_{2}}}$ is already an isometry. Moreover, one can use the appearance of $P_{H^{2}_{2}}$ in the $U_{\alpha}^{1}$ formula to verify that
$\text{Ran}(S_{\phi}^{1}|_{K_{\phi_{2}}})=z_{1}K_{\phi_{2}}\perp\text{Ran}(P_{\phi}P_{H^{2}_{2}}M_{\overline{\vartheta^{1}_{\alpha}}})=\text{Ran}(U_{\alpha}^{1}-S_{\phi}^{1}).$
This implies that $(U_{\alpha}^{1}-S_{\phi}^{1})|_{K_{\phi_{2}}}\equiv 0.$ Thus, fully understanding $U_{\alpha}^{1}$ would require us to more fully understand
@@
$=P_{\phi}P_{H^{2}_{2}}\left(\alpha\frac{\overline{(B_{1}\phi)(z)}}{1-\alpha\overline{\phi(0,z_{2})}}f\right)$
@@
Because of the involvement of $P_{H^{2}_{2}}$ and $g(z_{2})$, this formula does not appear to nicely mimic the one-variable situation. Still, for some special functions, we can compute parts of this formula more directly. For example, if we set
@@
If we expand this as a power series in $L^{2},$ we can see that there are no terms with positive $z_{2}$ power. That means that the projection of this function onto $H^{2}_{2}$ must be a constant function. One can compute this directly as:
$\left\langle\frac{\overline{(B_{1}\phi_{1})(z_{1})}(B_{1}\phi_{1})(z_{1})}{1-\alpha\overline{\phi_{1}(0)\phi_{2}(z_{2})}},1\right\rangle_{L^{2}}$
$=\left\langle\overline{(B_{1}\phi_{1})(z_{1})}(B_{1}\phi_{1})(z_{1}),1\right\rangle_{L^{2}}\left\langle\frac{1}{1-\alpha\overline{\phi_{1}(0)\phi_{2}(z_{2})}},1\right\rangle_{L^{2}}$
$=\|B_{1}\phi_{1}\|^{2}_{L^{2}}\frac{1}{1-\alpha\overline{\phi_{1}(0)\phi_{2}(0)}}\neq 0.$
@@
$P_{\phi}P_{H^{2}_{2}}\overline{\vartheta^{1}_{\alpha}}F=P_{\phi}\left(\|B_{1}\phi_{1}\|^{2}_{L^{2}}\frac{1}{1-\alpha\overline{\phi_{1}(0)\phi_{2}(0)}}\right).$
@@
6.2. A Singular Rational Inner Function
@@
$\phi(z)=\frac{2z_{1}z_{2}-z_{1}-z_{2}}{2-z_{1}-z_{2}}$ (6.1)
which has a singularity at $\zeta=(1,1)$. Note that $\phi(1,z_{2})=\phi(z_{1},1)=-1$ so that $\alpha=-1$ is an exceptional value of this RIF. In fact, since the level sets of $\phi$ can be parameterized as
@@
this is the sole exceptional value of this function. The Clark measures of $\phi$ are described for $\alpha\neq-1$ by
@@
and by Theorem 3.4, the Clark embedding operator $J_{\alpha}\colon K_{\phi}\to L^{2}(\sigma_{\alpha})$ is unitary. Note that while $\bigcap_{\alpha\in\mathbb{T}}\mathrm{supp}\,\sigma_{\alpha}=\{(1,1)\}$, we have $\sigma_{\alpha}(\{1,1\})=0$ for each $\alpha$, so there is no contradiction to mutual singularity of the Clark measures. See [9, Example 5.1] for further details including a discussion of what happens in the exceptional case $\alpha=-1$.
Figure 2. Support sets for $\sigma_{\alpha}$ for $\phi=\frac{2z_{1}z_{2}-z_{1}-z_{2}}{2-z_{1}-z_{2}}$ and $\alpha=1$ (black), $\alpha=e^{i\frac{\pi}{4}}$ (gray), $\alpha=e^{i\frac{\pi}{2}}$ (orange), $\alpha=e^{-i\frac{\pi}{2}}$ (pink), and the exceptional value $\alpha=-1$ (red). Singular point $(1,1)\simeq(0,0)$ marked in red.
@@
which is no longer a bounded rational function in $\mathbb{D}^{2}$, viz. [27, p. 1265] (or see [10, Theorem 1.2] for a more general explanation).

### v2

@@ 5. RIF Level Sets and Taylor Joint Spectrum
First, we show that on $L^{2}(\sigma_{\alpha})$, we have $\boldsymbol{\sigma}(M_{z_{1}},M_{z_{2}})=\mathcal{C}_{\alpha}$. By Theorem 3.3, the set $\mathcal{C}_{\alpha}$ is the support of $\sigma_{\alpha}$. To apply Theorem 5.1, note that $\mathcal{C}_{\alpha}$ is the closure of the set
@@
as needed. Then $\boldsymbol{\sigma}(M_{z_{1}},M_{z_{2}})=\mathcal{C}_{\alpha}$ follows immediately from Theorem 5.1.
To move this result to the model space $K_{\phi}$, note that Corollary 4.6 implies that $J_{\alpha}$ is unitary and there are unitary maps $U_{\alpha}^{1},U_{\alpha}^{2}:K_{\phi}\rightarrow K_{\phi}$ such that $U^{j}_{\alpha}=J_{\alpha}^{*}M_{z_{j}}J_{\alpha}$ for $j=1,2$. Since $M_{z_{1}},M_{z_{2}}$ commute on $L^{2}(\sigma_{\alpha})$, the operators $U_{\alpha}^{1},U_{\alpha}^{2}$ commute on $K_{\phi}$. Moreover,
@@
where $\vartheta^{2}_{\alpha}(z)=\bar{\alpha}\frac{(B_{2}\phi)(z)}{1-\bar{\alpha}\phi(z_{1},0)}$ and $B_{2}$ denotes the backward shift in $z_{2}$. (As an aside, while Corollary 4.6 only states the information about $U_{\alpha}^{1}$, the proof immediately implies the corresponding information about $U_{\alpha}^{2}$.) Thus, the operators $U^{1}_{\alpha},U^{2}_{\alpha}$ are commuting unitary perturbations of the compressed shifts $S_{\phi}^{1},S_{\phi}^{2}$. By above, $\boldsymbol{\sigma}(M_{z_{1}},M_{z_{2}})=\mathcal{C}_{\alpha}$. Since the Taylor joint spectrum is invariant under unitary conjugation, this gives $\boldsymbol{\sigma}(U_{\alpha}^{1},U_{\alpha}^{2})=\mathcal{C}_{\alpha}$. ∎
@@
There are several techniques for constructing rational inner functions $\phi$ whose level sets $\mathcal{C}_{\alpha}$ have prescribed properties, see for instance [17, Section 7]. By Theorem 5.4, this in turn allows us to prescribe such curve subsets of $\mathbb{T}^{2}$ as Taylor spectra of natural pairs of commuting unitaries on model spaces.
@@ 6.1. Finite Blaschke Products
In this subsection, we use inner functions constructed from finite Blaschke products to explore our various results. Specifically, let $\phi_{1}$ and $\phi_{2}$ be finite Blaschke products (with at least one nonconstant) and consider $\phi(z)=\phi(z_{1})\phi(z_{2})$.
@@
First, let us assume that one of $\phi_{1}$ or $\phi_{2}$ is constant, so that $\phi$ is a function of one variable. Then for each $\alpha\in\mathbb{T},$ the level set $\mathcal{C}_{\alpha}$ will contain lines of either the form $\mathbb{T}\times\{\tau\}$ or $\{\tau\}\times\mathbb{T}$ for some $\tau\in\mathbb{T}.$ Then basically by applying the argument in the proof of Proposition 3.10 in [14], one can show that the isometry $J_{\alpha}$ linking $K_{\phi}$ and $L^{2}(\sigma_{\alpha})$ is not unitary. Thus, our machinery breaks down and we cannot apply Theorem 4.3 to obtain a unitary perturbation of $S_{\phi}^{1}.$
@@
For this reason, it makes sense to restrict attention to $\phi_{1},\phi_{2}$ that are both nonconstant. In this situation, since $\phi$ also has no boundary singularities, every $\alpha\in\mathbb{T}$ must be generic for $\phi$. From this, we can apply Corollary 4.6 to obtain the formula
@@
for the associated perturbation of $S_{\phi}^{1}$ to the unitary $U_{\alpha}^{1}$.
@@
If $\phi_{1}(0)=0$, then $\psi_{1}(z_{1}):=\phi_{1}(z_{1})/z_{1}$ is also a finite Blaschke product and $\phi(z)=z_{1}\psi_{1}(z_{1})\phi_{2}(z_{2})$ with $\phi(0,z_{2})\equiv 0$. In this case, we have a particularly nice formula for $U_{\alpha}^{1}.$ We can apply Corollary 4.5 to conclude that
@@
where $K_{\phi_{2}}$ is invariant under $T_{1},$ the shift in $z_{1}$ and thus $K_{\phi_{2}}\subseteq S_{1}^{\mathrm{max}}$. By Remark 4.8, $\phi$ satisfies the assumptions of Theorem 4.9 and so, $(U_{\alpha}^{1}-S_{\phi}^{1})|_{K_{\phi_{2}}}\equiv 0.$ Thus, fully understanding $U_{\alpha}^{1}$ would require us to more fully understand
@@
$=P_{\phi}P_{H^{2}_{2}}\left(\alpha\frac{\overline{(B_{1}\phi)(z)}}{1-\alpha\overline{\phi(0,z_{2})}}f(z)\right)$
@@
Because of the involvement of $P_{H^{2}_{2}}$ and $g(z_{2})$, this formula does not appear to nicely mimic the one-variable situation. Still, for some special functions, we can compute parts of this formula more directly. For example, if we set $h=B_{1}\phi_{1}$ and $g=1$ to get
@@
If we expand this as a Fourier series in $L^{2},$ we can see that there are no terms with positive $z_{2}$ power. That means that the projection of this function onto $H^{2}_{2}$ must be a constant function. One can compute this directly as:
$\alpha\left\langle\frac{\overline{(B_{1}\phi_{1})(z_{1})}(B_{1}\phi_{1})(z_{1})}{1-\alpha\overline{\phi_{1}(0)\phi_{2}(z_{2})}},1\right\rangle_{L^{2}}$
$=\alpha\left\langle\overline{(B_{1}\phi_{1})(z_{1})}(B_{1}\phi_{1})(z_{1}),1\right\rangle_{L^{2}}\left\langle\frac{1}{1-\alpha\overline{\phi_{1}(0)\phi_{2}(z_{2})}},1\right\rangle_{L^{2}}$
$=\|B_{1}\phi_{1}\|^{2}_{L^{2}}\frac{\alpha}{1-\alpha\overline{\phi_{1}(0)\phi_{2}(0)}}\neq 0.$
@@
$P_{\phi}P_{H^{2}_{2}}\overline{\vartheta^{1}_{\alpha}}F=P_{\phi}\left(\|B_{1}\phi_{1}\|^{2}_{L^{2}}\frac{\alpha}{1-\alpha\overline{\phi_{1}(0)\phi_{2}(0)}}\right).$
@@
6.2. A Rational inner function
@@
(6.1) $\phi(z)=\frac{2z_{1}z_{2}-z_{1}-z_{2}}{2-z_{1}-z_{2}},$
which has a singularity at $(1,1)$ and often serves as an important example in RIF theory (e.g. [33, Section 3], [16, Example 1], and [14, Example 5.1]).
Note that $\phi(1,z_{2})=\phi(z_{1},1)=-1$ so that $\alpha=-1$ is an exceptional value of this RIF. In fact, since the level sets of $\phi$ for other $\alpha$ can be parameterized as
@@
this is the sole exceptional value of this function. A collection of level sets is graphed in Figure 2. The Clark measures of $\phi$ are described for $\alpha\neq-1$ by
@@
and by Theorem 3.4, the Clark embedding operator $J_{\alpha}\colon K_{\phi}\to L^{2}(\sigma_{\alpha})$ is unitary. Note that while $\bigcap_{\alpha\in\mathbb{T}}\mathrm{supp}\,\sigma_{\alpha}=\{(1,1)\}$, we have $\sigma_{\alpha}(\{1,1\})=0$ for each $\alpha$, so there is no contradiction to mutual singularity of the Clark measures.
Figure 2. Support sets in $\mathbb{T}^{2}\simeq[-\pi,\pi)^{2}$ for the Clark measures $\sigma_{\alpha}$ for $\phi=\frac{2z_{1}z_{2}-z_{1}-z_{2}}{2-z_{1}-z_{2}}$ and $\alpha=1$ (black), $\alpha=e^{i\frac{\pi}{4}}$ (gray), $\alpha=e^{i\frac{\pi}{2}}$ (orange), $\alpha=e^{-i\frac{\pi}{2}}$ (pink), and the exceptional value $\alpha=-1$ (red). Singular point $(1,1)\simeq(0,0)$ marked in red.
@@
which is no longer a bounded rational function in $\mathbb{D}^{2}$, viz. [33, p. 1265] (or see [15, Theorem 1.2] for a more general explanation). See [14, Example 5.1] for further details including a discussion of what happens in the exceptional case $\alpha=-1$.
The value $\alpha=1$ is generic, and we have
$\vartheta^{1}_{1}(z)=-\frac{(z_{2}-1)^{2}}{2-z_{1}-z_{2}}.$
The structure of the corresponding Clark measure also simplifies, and we have
$\int_{\mathbb{T}^{2}}fd\sigma_{1}=\int_{\mathbb{T}}f(\zeta_{1},\bar{\zeta}_{1})\frac{1}{2}|1-\zeta_{1}|^{2}dm(\zeta_{1}).$
Since $\phi(0,z_{2})\neq 0$ and $\phi(z_{1},0)\neq 0$, Proposition 4.7 shows that a projection is needed in the formulas for the unitary operators $U_{1}^{1},U_{1}^{2}$. By appealing to Agler decompositions, which were discussed in Subsection 3.3, we can nevertheless work out a more concrete representation for the Clark unitaries for $\alpha=1$.
First, we note that since $\deg\phi=(1,1)$ and $\phi$ has a singularity at $(1,1),$ Lemma 4.1 in [12] implies that $\phi$ has a unique Agler decomposition given by
(6.2) $1-\overline{\phi(w)}\phi(z)=(1-\bar{w}_{1}z_{1})g(z)\overline{g(w)}+(1-\bar{w}_{2}z_{2})h(z)\overline{h(w)},$
where
$g(z)=\frac{\sqrt{2}(1-z_{2})}{2-z_{1}-z_{2}}\quad\textrm{and}\quad h(z)=\frac{\sqrt{2}(1-z_{1})}{2-z_{1}-z_{2}}.$
Now set
$K_{1}(z,w)=h(z)\overline{h(w)}\ \ \text{ and }\ \ K_{2}(z,w)=g(z)\overline{g(w)}.$
By the uniqueness of the Agler kernels for this $\phi$, it must be the case that $\mathcal{H}(K_{1}),\mathcal{H}(K_{2})$ are closed subspaces of $H^{2}$ (where the reproducing kernel norm agrees with the $H^{2}$ norm) and thus, $\{h\}$ is an orthonormal basis for $\mathcal{H}(K_{1})$ and $\{g\}$ is an orthonormal basis for $\mathcal{H}(K_{2})$. Meanwhile, as mentioned in Subsection 3.3, we have the orthogonal decomposition
$S_{1}^{\mathrm{max}}=\bigoplus_{k=0}^{\infty}z_{1}^{k}\mathcal{H}(K_{1})\quad\textrm{and}\quad S_{2}^{\mathrm{min}}=\bigoplus_{l=0}^{\infty}z_{2}^{l}\mathcal{H}(K_{2}).$
This implies that $\{z_{1}^{k}h\}_{k=0}^{\infty}$ and $\{z_{2}^{l}g\}_{l=0}^{\infty}$ are orthonormal bases for $S_{1}^{\mathrm{max}}$ and $S_{2}^{\mathrm{min}}$ respectively. We also note that
$\vartheta_{1}^{1}(z)=-\frac{1-z_{2}}{\sqrt{2}}g(z).$
We can now determine the action of $U^{1}_{1}$ on the orthonormal basis of $S_{2}^{\mathrm{min}}$. First, we compute, using $\|g\|_{H^{2}}=1$ and $g\perp z_{2}^{k}g$, that for $w\in\mathbb{D}^{2},$
$\left(P_{H^{2}_{2}}M_{\overline{\vartheta_{1}^{1}}}gz_{2}^{n}\right)(w)$ $=\left\langle\overline{\vartheta_{1}^{1}(z)}g(z)z_{2}^{n},\frac{1}{1-\bar{w}_{2}z_{2}}\right\rangle_{L^{2}}$
$=\left\langle g(z)z_{2}^{n},\frac{\vartheta_{1}^{1}(z)}{1-\bar{w}_{2}z_{2}}\right\rangle_{H^{2}}$
$=\sum_{k=0}^{\infty}\frac{1}{\sqrt{2}}\langle g(z)z_{2}^{n},(z_{2}-1)g(z)z_{2}^{k}\rangle_{H^{2}}\,w_{2}^{k}$
$=\frac{1}{\sqrt{2}}\sum_{k=0}^{\infty}(\langle gz_{2}^{n-k-1},g\rangle_{H^{2}}-\langle gz_{2}^{n-k},g,\rangle_{H^{2}})w_{2}^{k}$
(6.5) $=\left\{\begin{array}[]{cc}-\frac{1}{\sqrt{2}}\,,&n=0,\\ \frac{1}{\sqrt{2}}(w_{2}^{n-1}-w_{2}^{n}),&n>0.\end{array}\right$.
From this we see that $\mathrm{Ran}(P_{H_{2}^{2}}M_{\overline{\vartheta_{1}^{1}}})\mid_{S_{2}^{\mathrm{min}}}=H_{2}^{2}$. It remains to apply the model space projection $P_{\phi}$ to (6.5). In order to do this, we first use (6.2) to get
$P_{\phi}(z_{2}^{k})(w)$ $=\left\langle z_{2}^{k},k_{w}^{\phi}\right\rangle_{H^{2}}$
$=\left\langle z_{2}^{k},\frac{1-\overline{\phi(w)}\phi(z)}{(1-\bar{w}_{1}z_{1})(1-\bar{w}_{2}z_{2})}\right\rangle_{H^{2}}$
$=\left\langle z_{2}^{k},\frac{\overline{g(w)}g(z)}{1-\bar{w}_{2}z_{2}}+\frac{\overline{h(w)}h(z)}{1-\bar{w}_{1}z_{1}}\right\rangle_{H^{2}}$
$=\sum_{l=0}^{k}\langle z_{2}^{k},g(z)z_{2}^{l}\rangle_{H^{2}}w_{2}^{l}g(w)+\langle z_{2}^{k},h\rangle_{H^{2}}h(w);$
note that the second inner product reduces to one term since powers of $z_{1}$ and $z_{2}$ are orthogonal. We now obtain
$(P_{\phi}P_{H^{2}_{2}}M_{\overline{\vartheta_{1}^{2}}}g)(w)$ $=P_{\phi}\left(-\frac{1}{\sqrt{2}}\right)$
$=-\frac{1}{\sqrt{2}}\left(\langle 1,g\rangle_{H^{2}}g(w)+\langle 1,h\rangle_{H^{2}}h(w)\right)$
$=-\frac{1}{\sqrt{2}}(\overline{g(0)}g(w)+\overline{h(0)}h(w))$
and for $n\geq 1$,
$(P_{\phi}P_{H^{2}_{2}}M_{\overline{\vartheta_{1}^{2}}}gz_{2}^{n})(w)$ $=\frac{1}{\sqrt{2}}P_{\phi}\left(z_{2}^{n-1}-z_{2}^{n}\right)$
$=\frac{1}{\sqrt{2}}\left(\sum_{l=0}^{n-1}\langle z_{2}^{n-1-l}-z_{2}^{n-l},g\rangle_{H^{2}}w_{2}^{l}-\overline{g(0)}w_{2}^{n}\right)g(w)$
$\quad+\frac{1}{\sqrt{2}}\langle z_{2}^{n-1}-z_{2}^{n},h\rangle_{H^{2}}h(w).$
On the other hand, $P_{\phi}P_{H^{2}_{2}}M_{\overline{\vartheta_{1}^{2}}}|_{S_{1}^{\mathrm{max}}}=0$ as guaranteed by Theorem 4.9.
As this relatively simple example illustrates, it does not seem to be easy to find a more concrete representation of Clark unitaries in the case when the underlying inner function has $\phi(0,z_{2})\neq 0$ and $\phi(z_{1},0)\neq 0$.